A company's invested capital is $13,000,000 and management has determined that the target rate of return on investment is 10%. Last year, the company produced 131,313 units and this year expects to units sales to be 10% above last year. The cost of the product is estimated to be $13 per unit. What is the target operating income per unit? (Round any intermediary calculations to the nearest unit and your final answer to the nearest cent.)

A) $9.00

B) $9.90

C) $13.00

D) $6.50
Answer #1

Target return on investment = 13,000,000*10%

= 1,300,000

Expected sales = 131,313 + 10%

= 144,444 units

Target operating income per unit = Target return/Expected sales

= 1,300,000/144,444

= 9

Option A is the answer